# Beyond Correlation: Causal Multi-View Unsupervised Feature Selection Learning

### v1

@@ Abstract
Index Terms:
Unsupervised Feature Selection, Multi-view Unlabeled Data, Mitigating Spurious Correlations.
@@ 1 Introduction
With the widespread availability of high-dimensional, unlabeled multi-view data in practical applications, multi-view unsupervised feature selection (MUFS) has attracted increasing attention in recent years [1, 2, 3, 4]. It aims to identify representative features from multiple heterogeneous feature sets, thus effectively addressing the “curse of dimensionality” problem and improving the performance of downstream tasks. Existing MUFS methods typically leverage the cross-view local structures to learn clustering labels. Simultaneously, they capture correlations between features and the inferred clustering labels to guide feature selection, which has shown promising performance in practice [5, 6].
Generally, the key question is whether the correlations captured by existing MUFS methods are sufficiently reliable to guide feature selection. If not, what factors undermine this reliability, and how can it be improved? Investigating this question can deepen our understanding of the MUFS problem, reveal the limitations of existing methods, and inspire more effective solutions. The causal relationships are widely acknowledged to be among the most reliable and interpretable foundations for downstream tasks [7]. For example, in animal image clustering, causal features (e.g., facial structure) inherently distinguish dogs from other animals, and selecting these features can therefore improve clustering performance. In contrast, non-causal features (e.g., background elements) do not provide such discriminative information, and their inclusion may degrade clustering performance. This observation motivates us to analyze MUFS from a causal perspective. To this end, we introduce a novel structural causal model (SCM) in Section 2. The SCM demonstrates that confounding factors that influence both features and clustering labels can induce spurious correlations. For instance, background features may be spuriously correlated with clustering labels if the dataset is biased towards specific environments (e.g., cats are often photographed indoors, while dogs are frequently photographed on grass). Such distribution bias may induce potential confounding, leading to spurious correlations between grass features and clustering labels [8, 9]. Since existing MUFS methods ignore such confounding effects, they often rely on spurious correlations to select non-causal features, such as grass features in the aforementioned example.
@@
Recognizing this limitation naturally raises the question of how to mitigate such spurious correlations and identify causally informative features. Addressing this challenge in unsupervised settings is particularly difficult due to the lack of label information. Although several causal feature selection methods have been proposed recently, their reliance on label information makes them inapplicable to unsupervised scenarios [10, 11, 12]. Based on Pearl’s back-door criterion, spurious correlations can be mitigated through confounder adjustment [13]. Accordingly, several confounder balancing methods have been proposed in the causal literature, which construct sample weights to balance confounder distributions, thereby blocking backdoor paths and reducing spurious correlations [14, 15]. However, most existing confounder balancing methods treat all observed features as confounders, neglecting non-confounding features that only affect the clustering labels or the treatment features. Balancing these non-confounders may result in inaccurate causality identification by existing methods [16]. Therefore, how to adaptively separate confounding factors and mitigate spurious correlations in order to identify causally informative features from unlabeled multi-view data remains an urgent challenge in practical applications.
@@ 3.1 Formulation of CAUSA
where $\bm{W}^{(v)}\in\mathbb{R}^{d_{v}\times c}$ denotes the feature selection matrix of the $v$-th view, $\bm{F}\in\mathbb{R}^{n\times c}$ represents the consensus clustering label across different views, and $c$ is the dimension of the low-dimensional space. Moreover, $\bm{L}$ denotes the cross-view Laplacian matrix, constructed using the method proposed in [17]. In Eq. (2), the $i$-th row of $\bm{W}^{(v)}$ captures the dependence between the feature $\bm{X}_{i\cdot}^{(v)}$ and the clustering label $\bm{F}$, where a stronger dependence indicates a greater contribution of $\bm{X}_{i\cdot}^{(v)}$ to distinguishing clusters. Accordingly, $\|\bm{W}_{i\cdot}^{(v)}\|_{2}^{2}$ can serve as a quantitative measure of the importance of the $i$-th feature.
@@
In Theorem 1, we demonstrate that spurious correlations between non-causal features and clustering labels can be eliminated by ensuring that the confounder distribution remains consistent across different values of the non-causal features. This is consistent with Pearl’s backdoor criterion [13], which asserts that adjusting for confounders (i.e., ensuring $\mathcal{C}^{(v)}\perp\mathcal{U}^{(v)}$ in the adjusted distribution) blocks spurious paths and enables causal identification. Motivated by this observation, we propose a novel causal regularization module that learns sample weights to balance the distribution of confounders, which are adaptively identified from multiple views.
Specifically, given a treatment feature $\bm{X}_{r\cdot}^{(v)}$ in view $v$, we consider the remaining features that are associated with both $\bm{X}_{r\cdot}^{(v)}$ and the clustering label as confounders, denoted by $\bm{C}^{r^{(v)}}$. We identify these confounders using our proposed adaptive confounder separation method, which will be introduced later. Then, the samples characterized by confounders $\bm{C}^{r^{(v)}}$ are partitioned into the treatment and control groups according to values of the treatment feature $\bm{X}_{r\cdot}^{(v)}$. Without loss of generality, we define the treatment feature as a binary variable taking values of 0 or 1, assigning samples to the treatment group when the value is 1 and to the control group when it is 0. To mitigate confounding bias and capture true causality of $\bm{X}_{r\cdot}^{(v)}$, we learn a set of sample weights to balance the confounder distribution between treatment and control groups. Meanwhile, we employ Maximum Mean Discrepancy (MMD) [18] to quantify distributional differences, as this kernel-based metric characterizes complex nonlinear relationships by embedding distributions into a Reproducing Kernel Hilbert Space (RKHS). The corresponding objective function is formulated below:
@@
where $\bm{\tau}^{(v)}\in\mathbb{R}^{n\times 1}$ denotes the vector of sample weights in the $v$-th view. The sets $\nabla_{r}^{(v)}$ and $\Delta_{r}^{(v)}$ indicate the indices of samples in the treatment and control groups, respectively, with $|\nabla_{r}^{(v)}|$ denoting the cardinality of $\nabla_{r}^{(v)}$. Moreover, $\phi(\tau_{i}^{(v)}\!\bm{C}_{\cdot i}^{r^{(v)}})=\kappa(\tau_{i}^{(v)}\!\bm{C}_{\cdot i}^{r^{(v)}},\cdot)$ represents the feature map induced by the kernel $\kappa(\cdot,\cdot)$, and $\|\cdot\|_{\mathcal{H}_{\kappa}}$ denotes the RKHS norm [19].
To enhance the balancing process in Eq. (4), we propose an adaptive confounder separation method that, unlike conventional causal approaches, does not treat all observed features as potential confounders. Specifically, for the treatment feature $\bm{X}_{r\cdot}^{(v)}$, the remaining features, denoted as $\bm{Z}^{r^{(v)}}\!\!\!=\!\bm{X}^{(v)}\backslash\bm{X}_{r\cdot}^{(v)}\!\in\!\mathbb{R}^{(d_{v}-1)\times n}$, can typically be divided into three subsets: confounders, adjustment variables, and instrumental variables [20]. Due to the lack of prior knowledge about the true causal relationships, we propose a data-driven approach to identify confounders from $\bm{Z}^{r^{(v)}}$ by exploiting their dual associations with both treatment feature and clustering label. More specifically, we introduce a learnable confounding indicator vector $\bm{e}^{r^{(v)}}\in\{0,1\}^{(d_{v}-1)\times 1}$, where ${e}_{i}^{r^{(v)}}=1$ if the $i$-th feature of $\bm{Z}^{r^{(v)}}$ is a confounder, and ${e}_{i}^{r^{(v)}}=0$ otherwise. Thus, $\bm{C}^{r^{(v)}}$ can be written as $\bm{C}^{r^{(v)}}=(\bm{e}^{r^{(v)}}\textbf{1}^{\top})\odot\bm{Z}^{r^{(v)}}$, where $\odot$ denotes the Hadamard product. By simultaneously maximizing the associations between $\bm{C}^{r^{(v)}}$ and both the clustering labels and the treatment feature, the indicator vector $\bm{e}^{r^{(v)}}$ can be adaptively optimized to extract the proxy confounders from $\bm{Z}^{r^{(v)}}$. Accordingly, Eq. (4) can be reformulated as follows:
@@
Furthermore, to capture causal contributions of all features, a common strategy is to iteratively treat each view-specific feature as the treatment feature, while learning corresponding sample weights to mitigate confounding bias [21, 22]. However, this strategy becomes computationally prohibitive in high-dimensional settings. Given that different views share the same set of samples, we propose learning view-shared sample weights to achieve confounding balance. Moreover, prior studies have demonstrated that mitigating confounding bias for a carefully selected subset of prototype features can help reduce spurious correlations across the entire feature set [23, 24]. Accordingly, we propose balancing confounders for $m~(m\ll d_{v})$ causally representative prototype features within each view, rather than all view-specific features. To this end, we introduce a causality-guided hierarchical feature clustering method to select $m$ prototype features from each view, with their indices denoted as $\Lambda^{(v)}$. The inter-feature distances are defined by the differences in their causal contributions as represented in $\bm{W}^{(v)}$, which is jointly learned during the confounder balancing process. We then learn view-shared sample weights to balance confounding distributions associated with the $m$ prototype features and formulate the proposed causal regularization module as follows:
@@ 4 Optimization and Analyses
where $\bm{A}^{(v)}=\alpha\bm{X}^{(v)}\bm{\Gamma}\bm{X}^{(v)\top}+\lambda\bm{D}^{(v)}$, $\bm{\Gamma}=\operatorname{diag}(\bm{\tau})$, $\bm{D}^{(v)}$ is a diagonal matrix whose $i$-th diagonal entry is given by ${D}_{ii}^{(v)}=1/2\sqrt{\|\bm{W}_{i\cdot}^{(v)}\|_{2}^{2}+\epsilon}$ ($\epsilon$ is a small constant to prevent the denominator from vanishing), and $\bm{B}^{(v)}=\alpha\bm{X}^{(v)}\bm{\Gamma}\bm{F}$. Problem (8) can be efficiently solved using the generalized power iteration (GPI) algorithm [25].
Update $\bm{F}$ by Fixing Others. With other variables fixed, we can formulate the Lagrangian function w.r.t. $\bm{F}$, and obtain the following optimal solution of $\bm{F}$ by applying the Karush–Kuhn–Tucker (KKT) conditions [26]:
@@
Update $\bm{E}^{(v)}$ by Fixing Others. Following commonly adopted relaxation strategies for discrete constraints [27], we relax $\bm{E}^{(v)}$ to take continuous values in the range $[0,1]$, thereby reformulating the optimization problem w.r.t. $\bm{E}^{(v)}$ as follows:
@@
where $\varrho$ is a large constant used to enforce sparsity of $\bm{E}_{\cdot r}^{(v)}$. Problem (10) can be efficiently solved using the proximal gradient descent method [28].
Update $\bm{\tau}$ by Fixing Others. Based on the empirical estimation of MMD from [29], the optimization problem w.r.t. $\bm{\tau}$ can be reformulated as follows:
@@
where $H_{ij}\!=\!\sum_{v=1}^{V}\sum_{r=1}^{m}K_{ij}^{r^{(v)}}\bm{C}_{\cdot i}^{r^{(v)}\top}\!\bm{C}_{\cdot j}^{r^{(v)}}$, $K_{ij}^{r^{(v)}}\!=\!\Omega_{ri}^{(v)}\Omega_{rj}^{(v)}\!-\!2\Omega_{ri}^{(v)}(1\!-\!\Omega_{rj}^{(v)})+(1\!-\!\Omega_{ri}^{(v)})(1\!-\!\Omega_{rj}^{(v)})$, and $\Omega_{ri}^{(v)}$ denotes an indicator variable for view $v$ which equals 1 if the $i$-th sample is assigned to the treatment group based on the value of the $r$-th treatment feature, and 0 otherwise. Moreover, $\bm{g}$ is a vector whose $i$-th entry is given by $g_{i}=\sum_{v=1}^{V}\|\bm{X}_{\cdot i}^{(v)\top}\bm{W}^{(v)}-\bm{F}_{i\cdot}\|_{2}$, where $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm. Problem (11) is a standard simplex projection problem, which can be efficiently solved using the method proposed in [30].
@@ 5.1 Experimental Settings
Datasets. We conduct experiments on six real-world multi-view datasets, including MSRA [31], MSRC [32], Prokaryotic [33], SensIT [34], CCV [35], and FMNIST [36]. Table I summarizes the statistical details of these datasets.
Compared Methods. We compare the proposed CAUSA with several SOTA methods, including SPDFS [37], BDGFS [38], GCDUFS [39], UKMFS [6], WLTL [40], JMVFG [41], OEDFS [42], CvLP-DCL [5], as well as a baseline (AllFea) that utilizes all original features.
Comparison Schemes. To ensure a fair comparison, we employ the grid search strategy to tune the parameters of all methods and report their respective optimal results. Simultaneously, the parameters $\alpha$, $\beta$ and $\lambda$ of our method are tuned over the range $\{10^{-3},10^{-2},10^{-1},1,10^{1},10^{2},10^{3}\}$. Because it is challenging to determine the optimal number of features [43], we vary the feature selection ratios from 10% to 50% in 10% increments across all datasets. Following a common practice for assessing MUFS [44, 45], we employ two widely used clustering metrics, Clustering Accuracy (ACC) and Normalized Mutual Information (NMI), to evaluate the quality of selected features. We run K-means clustering 50 times on the selected features and report the average result. All experiments were conducted using MATLAB R2022b on a desktop equipped with an Intel Core i9-10900 CPU (2.80 GHz) and 64 GB of RAM.

### v2

@@ 1 Introduction
With the widespread availability of high-dimensional, unlabeled multi-view data in practical applications, multi-view unsupervised feature selection (MUFS) has attracted increasing attention in recent years [li2024exploring, zhang2024scalable, zhang2024efficient, huang2025time]. It aims to identify representative features from multiple heterogeneous feature sets, thus effectively addressing the “curse of dimensionality” problem and improving the performance of downstream tasks. Existing MUFS methods typically leverage the cross-view local structures to learn clustering labels. Simultaneously, they capture correlations between features and the inferred clustering labels to guide feature selection, which has shown promising performance in practice [CvLPDCL, UKMFS].
Generally, the key question is whether the correlations captured by existing MUFS methods are sufficiently reliable to guide feature selection. If not, what factors undermine this reliability, and how can it be improved? Investigating this question can deepen our understanding of the MUFS problem, reveal the limitations of existing methods, and inspire more effective solutions. The causal relationships are widely acknowledged to be among the most reliable and interpretable foundations for downstream tasks [feuerriegel2024causal]. For example, in animal image clustering, causal features (e.g., facial structure) inherently distinguish dogs from other animals, and selecting these features can therefore improve clustering performance. In contrast, non-causal features (e.g., background elements) do not provide such discriminative information, and their inclusion may degrade clustering performance. This observation motivates us to analyze MUFS from a causal perspective. To this end, we introduce a novel structural causal model (SCM) in Section 2. The SCM demonstrates that confounding factors that influence both features and clustering labels can induce spurious correlations. For instance, background features may be spuriously correlated with clustering labels if the dataset is biased towards specific environments (e.g., cats are often photographed indoors, while dogs are frequently photographed on grass). Such distribution bias may induce potential confounding, leading to spurious correlations between grass features and clustering labels [kuang2021balance, huang2025debiasing]. Since existing MUFS methods ignore such confounding effects, they often rely on spurious correlations to select non-causal features, such as grass features in the aforementioned example.
@@
Recognizing this limitation naturally raises the question of how to mitigate such spurious correlations and identify causally informative features. Addressing this challenge in unsupervised settings is particularly difficult due to the lack of label information. Although several causal feature selection methods have been proposed recently, their reliance on label information makes them inapplicable to unsupervised scenarios [ling2022light, quinzan2023drcfs, mangal2025ate]. Based on Pearl’s back-door criterion, spurious correlations can be mitigated through confounder adjustment [pearl2014confounding]. Accordingly, several confounder balancing methods have been proposed in the causal literature, which construct sample weights to balance confounder distributions, thereby blocking backdoor paths and reducing spurious correlations [shen2018causally, zubizarreta2015stable]. However, most existing confounder balancing methods treat all observed features as confounders, neglecting non-confounding features that only affect the clustering labels or the treatment features. Balancing these non-confounders may result in inaccurate causality identification by existing methods [liu2023debiased]. Therefore, how to adaptively separate confounding factors and mitigate spurious correlations in order to identify causally informative features from unlabeled multi-view data remains an urgent challenge in practical applications.
@@ 3.1 Formulation of CAUSA
where $\bm{W}^{(v)}\in\mathbb{R}^{d_{v}\times c}$ denotes the feature selection matrix of the $v$-th view, $\bm{F}\in\mathbb{R}^{n\times c}$ represents the consensus clustering label across different views, and $c$ is the dimension of the low-dimensional space. Moreover, $\bm{L}$ denotes the cross-view Laplacian matrix, constructed using the method proposed in [zhou2006learning]. In Eq. (2), the $i$-th row of $\bm{W}^{(v)}$ captures the dependence between the feature $\bm{X}_{i\cdot}^{(v)}$ and the clustering label $\bm{F}$, where a stronger dependence indicates a greater contribution of $\bm{X}_{i\cdot}^{(v)}$ to distinguishing clusters. Accordingly, $\|\bm{W}_{i\cdot}^{(v)}\|_{2}^{2}$ can serve as a quantitative measure of the importance of the $i$-th feature.
@@
In Theorem 1, we demonstrate that spurious correlations between non-causal features and clustering labels can be eliminated by ensuring that the confounder distribution remains consistent across different values of the non-causal features. This is consistent with Pearl’s backdoor criterion [pearl2014confounding], which asserts that adjusting for confounders (i.e., ensuring $\mathcal{C}^{(v)}\perp\mathcal{U}^{(v)}$ in the adjusted distribution) blocks spurious paths and enables causal identification. Motivated by this observation, we propose a novel causal regularization module that learns sample weights to balance the distribution of confounders, which are adaptively identified from multiple views.
Specifically, given a treatment feature $\bm{X}_{r\cdot}^{(v)}$ in view $v$, we consider the remaining features that are associated with both $\bm{X}_{r\cdot}^{(v)}$ and the clustering label as confounders, denoted by $\bm{C}^{r^{(v)}}$. We identify these confounders using our proposed adaptive confounder separation method, which will be introduced later. Then, the samples characterized by confounders $\bm{C}^{r^{(v)}}$ are partitioned into the treatment and control groups according to values of the treatment feature $\bm{X}_{r\cdot}^{(v)}$. Without loss of generality, we define the treatment feature as a binary variable taking values of 0 or 1, assigning samples to the treatment group when the value is 1 and to the control group when it is 0. To mitigate confounding bias and capture true causality of $\bm{X}_{r\cdot}^{(v)}$, we learn a set of sample weights to balance the confounder distribution between treatment and control groups. Meanwhile, we employ Maximum Mean Discrepancy (MMD) [gretton2006kernel] to quantify distributional differences, as this kernel-based metric characterizes complex nonlinear relationships by embedding distributions into a Reproducing Kernel Hilbert Space (RKHS). The corresponding objective function is formulated below:
@@
where $\bm{\tau}^{(v)}\in\mathbb{R}^{n\times 1}$ denotes the vector of sample weights in the $v$-th view. The sets $\nabla_{r}^{(v)}$ and $\Delta_{r}^{(v)}$ indicate the indices of samples in the treatment and control groups, respectively, with $|\nabla_{r}^{(v)}|$ denoting the cardinality of $\nabla_{r}^{(v)}$. Moreover, $\phi(\tau_{i}^{(v)}\!\bm{C}_{\cdot i}^{r^{(v)}})=\kappa(\tau_{i}^{(v)}\!\bm{C}_{\cdot i}^{r^{(v)}},\cdot)$ represents the feature map induced by the kernel $\kappa(\cdot,\cdot)$, and $\|\cdot\|_{\mathcal{H}_{\kappa}}$ denotes the RKHS norm [cui2020calibrated].
To enhance the balancing process in Eq. (4), we propose an adaptive confounder separation method that, unlike conventional causal approaches, does not treat all observed features as potential confounders. Specifically, for the treatment feature $\bm{X}_{r\cdot}^{(v)}$, the remaining features, denoted as $\bm{Z}^{r^{(v)}}\!\!\!=\!\bm{X}^{(v)}\backslash\bm{X}_{r\cdot}^{(v)}\!\in\!\mathbb{R}^{(d_{v}-1)\times n}$, can typically be divided into three subsets: confounders, adjustment variables, and instrumental variables [wu2022learning]. Due to the lack of prior knowledge about the true causal relationships, we propose a data-driven approach to identify confounders from $\bm{Z}^{r^{(v)}}$ by exploiting their dual associations with both treatment feature and clustering label. More specifically, we introduce a learnable confounding indicator vector $\bm{e}^{r^{(v)}}\in\{0,1\}^{(d_{v}-1)\times 1}$, where ${e}_{i}^{r^{(v)}}=1$ if the $i$-th feature of $\bm{Z}^{r^{(v)}}$ is a confounder, and ${e}_{i}^{r^{(v)}}=0$ otherwise. Thus, $\bm{C}^{r^{(v)}}$ can be written as $\bm{C}^{r^{(v)}}=(\bm{e}^{r^{(v)}}\textbf{1}^{\top})\odot\bm{Z}^{r^{(v)}}$, where $\odot$ denotes the Hadamard product. By simultaneously maximizing the associations between $\bm{C}^{r^{(v)}}$ and both the clustering labels and the treatment feature, the indicator vector $\bm{e}^{r^{(v)}}$ can be adaptively optimized to extract the proxy confounders from $\bm{Z}^{r^{(v)}}$. Accordingly, Eq. (4) can be reformulated as follows:
@@
Furthermore, to capture causal contributions of all features, a common strategy is to iteratively treat each view-specific feature as the treatment feature, while learning corresponding sample weights to mitigate confounding bias [ni2024feature, kuang2019treatment]. However, this strategy becomes computationally prohibitive in high-dimensional settings. Given that different views share the same set of samples, we propose learning view-shared sample weights to achieve confounding balance. Moreover, prior studies have demonstrated that mitigating confounding bias for a carefully selected subset of prototype features can help reduce spurious correlations across the entire feature set [wang2019blessings, vander2011new]. Accordingly, we propose balancing confounders for $m~(m\ll d_{v})$ causally representative prototype features within each view, rather than all view-specific features. To this end, we introduce a causality-guided hierarchical feature clustering method to select $m$ prototype features from each view, with their indices denoted as $\Lambda^{(v)}$. The inter-feature distances are defined by the differences in their causal contributions as represented in $\bm{W}^{(v)}$, which is jointly learned during the confounder balancing process. We then learn view-shared sample weights to balance confounding distributions associated with the $m$ prototype features and formulate the proposed causal regularization module as follows:
@@ 4 Optimization and Analyses
where $\bm{A}^{(v)}=\alpha\bm{X}^{(v)}\bm{\Gamma}\bm{X}^{(v)\top}+\lambda\bm{D}^{(v)}$, $\bm{\Gamma}=\operatorname{diag}(\bm{\tau})$, $\bm{D}^{(v)}$ is a diagonal matrix whose $i$-th diagonal entry is given by ${D}_{ii}^{(v)}=1/2\sqrt{\|\bm{W}_{i\cdot}^{(v)}\|_{2}^{2}+\epsilon}$ ($\epsilon$ is a small constant to prevent the denominator from vanishing), and $\bm{B}^{(v)}=\alpha\bm{X}^{(v)}\bm{\Gamma}\bm{F}$. Problem (8) can be efficiently solved using the generalized power iteration (GPI) algorithm [nie2017generalized].
Update $\bm{F}$ by Fixing Others. With other variables fixed, we can formulate the Lagrangian function w.r.t. $\bm{F}$, and obtain the following optimal solution of $\bm{F}$ by applying the Karush–Kuhn–Tucker (KKT) conditions [boyd2004convex]:
@@
Update $\bm{E}^{(v)}$ by Fixing Others. Following commonly adopted relaxation strategies for discrete constraints [xing2021fairness], we relax $\bm{E}^{(v)}$ to take continuous values in the range $[0,1]$, thereby reformulating the optimization problem w.r.t. $\bm{E}^{(v)}$ as follows:
@@
where $\varrho$ is a large constant used to enforce sparsity of $\bm{E}_{\cdot r}^{(v)}$. Problem (10) can be efficiently solved using the proximal gradient descent method [nitanda2014stochastic].
Update $\bm{\tau}$ by Fixing Others. Based on the empirical estimation of MMD from [kumagai2019unsupervised], the optimization problem w.r.t. $\bm{\tau}$ can be reformulated as follows:
@@
where $H_{ij}\!=\!\sum_{v=1}^{V}\sum_{r=1}^{m}K_{ij}^{r^{(v)}}\bm{C}_{\cdot i}^{r^{(v)}\top}\!\bm{C}_{\cdot j}^{r^{(v)}}$, $K_{ij}^{r^{(v)}}\!=\!\Omega_{ri}^{(v)}\Omega_{rj}^{(v)}\!-\!2\Omega_{ri}^{(v)}(1\!-\!\Omega_{rj}^{(v)})+(1\!-\!\Omega_{ri}^{(v)})(1\!-\!\Omega_{rj}^{(v)})$, and $\Omega_{ri}^{(v)}$ denotes an indicator variable for view $v$ which equals 1 if the $i$-th sample is assigned to the treatment group based on the value of the $r$-th treatment feature, and 0 otherwise. Moreover, $\bm{g}$ is a vector whose $i$-th entry is given by $g_{i}=\sum_{v=1}^{V}\|\bm{X}_{\cdot i}^{(v)\top}\bm{W}^{(v)}-\bm{F}_{i\cdot}\|_{2}$, where $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm. Problem (11) is a standard simplex projection problem, which can be efficiently solved using the method proposed in [duchi2008efficient].
@@ 5.1 Experimental Settings
Datasets. We conduct experiments on six real-world multi-view datasets, including MSRA [zhang2017latent], MSRC [huang2022multi], Prokaryotic [wen2023unpaired], SensIT [chang2011libsvm], CCV [wang2019multi], and FMNIST [xiao2017fashion]. Table I summarizes the statistical details of these datasets.
Compared Methods. We compare the proposed CAUSA with several SOTA methods, including SPDFS [SPDFS], BDGFS [BDGFS], GCDUFS [GCDUFS], UKMFS [UKMFS], WLTL [WLTL], JMVFG [JMVFG], OEDFS [OEDFS], CvLP-DCL [CvLPDCL], as well as a baseline (AllFea) that utilizes all original features.
Comparison Schemes. To ensure a fair comparison, we employ the grid search strategy to tune the parameters of all methods and report their respective optimal results. Simultaneously, the parameters $\alpha$, $\beta$ and $\lambda$ of our method are tuned over the range $\{10^{-3},10^{-2},10^{-1},1,10^{1},10^{2},10^{3}\}$. Because it is challenging to determine the optimal number of features [li2017feature], we vary the feature selection ratios from 10% to 50% in 10% increments across all datasets. Following a common practice for assessing MUFS [FSDK, hou2023adaptive], we employ two widely used clustering metrics, Clustering Accuracy (ACC) and Normalized Mutual Information (NMI), to evaluate the quality of selected features. We run K-means clustering 50 times on the selected features and report the average result. All experiments were conducted using MATLAB R2022b on a desktop equipped with an Intel Core i9-10900 CPU (2.80 GHz) and 64 GB of RAM.